### v2

# HausaNLP at SemEval-2025 Task 11: Hausa Text Emotion Detection

Abstract
This paper presents our approach to multi-label emotion detection in Hausa, a low-resource African language, for SemEval Track A. We fine-tuned AfriBERTa, a transformer-based model pre-trained on African languages, to classify Hausa text into six emotions: anger, disgust, fear, joy, sadness, and surprise. Our methodology involved data preprocessing, tokenization, and model fine-tuning using the Hugging Face Trainer API. The system achieved a validation accuracy of 74.00%, with an F1-score of 73.50%, demonstrating the effectiveness of transformer-based models for emotion detection in low-resource languages.
HausaNLP at SemEval-2025 Task 11: Hausa Text Emotion Detection
Sani Abdullahi Sani1,2, Salim Abubakar1,3, Falalu Ibrahim Lawan1,4, Abdulhamid Abubakar1, Maryam Bala1 1HausaNLP, 2University of the Witwatersrand, 3Federal Polytechnic Daura, 4Kaduna State University correspondence: saniabdullahisani1@students.wits.ac.za
1 Introduction
Emotion detection in text is a fundamental Natural Language Processing (NLP) task with far-reaching applications in sentiment analysis, social media monitoring, and mental health support. While significant progress has been made in high-resource languages, low-resource languages like Hausa remain underrepresented in the NLP landscape. A key challenge is the limited availability of annotated Hausa datasets for emotion detection, hindering the development of robust models. We address this gap by leveraging the newly available BRIGHTER dataset Muhammad et al. (2025a) and fine-tuning a transformer-based model for multi-label emotion detection in Hausa.
The scarcity of comprehensive emotion lexicons and annotated corpora complicates the development and evaluation of emotion detection systems for low-resource languages (Kabir et al., 2023; Al-Wesabi et al., 2023; Raychawdhary et al., 2023a). Furthermore, these languages exhibit rich morphological variations, syntax, and semantic differences that are not well-captured by models trained on high-resource languages (Marreddy et al., 2022; V R et al., 2023). Code-mixing, where multiple languages are used within the same text, further complicates emotion detection in linguistically diverse contexts (Raychawdhary et al., 2023a; Sonu et al., 2022).
We leverage AfriBERTa, a transformer model specifically trained on African languages and fine-tuned for multi-class emotion classification. We preprocessed and tokenized the text data using the AfriBERTa tokenizer, and then fine-tuned the model on the BRIGHTER Hausa dataset. This work contributes to emotion detection research in low-resource languages by demonstrating effective methods for adapting transformer models to Hausa. Our findings highlight the potential of leveraging pre-trained models like AfriBERTa (Ogueji et al., 2021) for emotion detection tasks in low-resource African languages.
2 Background
2.1 SemEval Task Overview
SemEval-2024 Track A, the Multi-label Emotion Detection task under Task 11: Bridging the Gap in Text-Based Emotion Detection (Muhammad et al., 2025b), focuses on classifying text from multiple languages into six emotion categories: anger, disgust, fear, joy, sadness, and surprise. While the task encompasses a wide range of languages, our team (HausaNLP) focused on Hausa, a Chadic language spoken across West and Central Africa by approximately 88 million people, including 54 million native speakers and 34 million second-language users primarily in Nigeria, Niger, and neighboring countries (Wolff, 2024; Eberhard et al., 2024). The system input is a Hausa text sample, and the output is a one-hot encoded vector indicating the presence or absence of each emotion. This task is particularly challenging due to the nuances of emotion expression in Hausa, as well as the limited availability of annotated data for low-resource languages.
2.2 Hausa Dataset
The Hausa dataset is divided into train (approximately 2,145 samples), validation (356 samples), and test (1,080 samples) splits. The complete distribution of the dataset is shown in Figure  1. Each sample is annotated with one or more emotion labels, represented as a one-hot encoded vector as shown in  1. The dataset exhibits class imbalance, with emotions like joy and sadness being more prevalent than fear and disgust. This imbalance posed a challenge during model training and evaluation.
3 Related Work on Low-Resource Emotion Detection
Researchers have proposed several approaches to address the challenges of emotion detection in low-resource settings. The AfriSenti-SemEval 2023 Shared Task provided annotated datasets for low-resource languages like Hausa and Igbo (Raychawdhary et al., 2023a). Techniques such as few-shot and zero-shot learning have been employed to augment datasets and improve model performance in data-scarce environments (Agarwal and Abbas, 2025; Hasan et al., 2024).
Pre-trained models like mBERT, Cross-Lingual Model - RoBERTa (XLM-R), and BanglaBERT have been fine-tuned for specific low-resource languages, showing improved performance in emotion detection tasks (Raychawdhary et al., 2023b; Kabir et al., 2023; Raychawdhary et al., 2024). Hybrid models combining lexicon features with transformers have also been explored to enhance emotion classification (Kabir et al., 2023). Deep learning models (e.g., Convolutional Neural Networks (CNNs), Deep Neural Networks (DNNs)) and hybrid approaches (e.g., TCSO-DGNN) have been proposed to better capture the emotional nuances in text (Jadon and Kumar, 2023; Bao and Su, ).
AfriBERTa, a transformer model pre-trained on a diverse corpus of African languages, has shown promising results for tasks like text classification and named entity recognition in Hausa and other African languages (Ogueji et al., 2021). Our work builds on these advancements by fine-tuning AfriBERTa for emotion detection in Hausa, addressing the unique challenges of low-resource language processing.
Future directions include multimodal emotion detection, incorporating additional data modalities (e.g., audio, video) to complement textual analysis (Sarbazi-Azad et al., 2021; Wang and Zhang, 2025), developing cross-lingual and multilingual models that generalize across multiple low-resource languages (Raychawdhary et al., 2023b, 2024), and addressing ethical considerations related to data privacy, bias, and deployment of emotion detection systems in diverse cultural contexts (Agarwal and Abbas, 2025).
Figure 1: Distribution of emotions in the Hausa text dataset. The bar plot shows the frequency of each emotion category (joy, sadness, anger, fear, surprise, and disgust) present in the dataset, obtained through manual annotation by three independent annotators
Table 1: Sample entries from the Hausa emotion detection dataset. The table displays five example entries from the dataset, including the Hausa text, its English translation, and the corresponding emotion labels assigned by human annotators.
ID Text Anger Disgust Fear Joy Sadness Surprise
1 Kotu Ta Yi Hukunci Kan Shari’ar Zaben Dan Majalisar PDP, Ta Yi Hukuncin Bazata 0 0 0 0 0 1
2 Toh fah inji ’yan magana suka ce “ana wata ga wata” 0 0 0 0 0 1
3 Bincike ya nuna yan Najeriya sun fi damuwa da rashin tsaro da talauci fiye da korona 0 0 1 0 1 0
4 Methodology
Figure 2: Multi-label emotion detection system overview. The diagram illustrates the complete multi-label emotion detection process, which includes data preprocessing, feature extraction using TF-IDF and word embeddings, classification with a BiLSTM network, and evaluation using F1-score
4.1 Data Preprocessing
To prepare the data for training, a preprocessing pipeline was implemented. First, the text data was cleaned by converting it to lowercase and stripping extra spaces. This step ensured consistency and improved tokenization. Next, we mapped the one-hot encoded labels to a single integer label representing the dominant emotion. This simplified the classification task into a multi-class problem with six classes.
4.2 Tokenization and Formatting
We tokenized the preprocessed text data using the AfriBERTa tokenizer, a pre-trained tokenizer designed for African languages, including Hausa. The tokenizer was configured to truncate sequences longer than 128 tokens and pad shorter sequences to ensure uniform input sizes. The tokenized dataset was then formatted into PyTorch tensors, containing the input_ids, attention_mask, and label columns. This prepared the data for the transformer model.
4.3 Model Selection and Fine-Tuning
We selected the AfriBERTa Small model, a compact version of the AfriBERTa architecture, primarily due to compute limitations. We fine-tuned the model for sequence classification with six output labels corresponding to the six emotions.
The fine-tuning process was conducted using the Hugging Face Trainer Application Programming Interface (API). The training arguments were configured with a learning rate of 2e-5, a batch size of 8, and 5 epochs. To optimize training, we enabled mixed precision (fp16) and applied a warmup schedule with 500 steps to gradually increase the learning rate at the start of training. The model was evaluated on the validation set after each epoch, and the best model was saved based on validation accuracy.
4.4 Evaluation Metrics
We evaluated the model’s performance during training using standard classification metrics: accuracy, precision, recall, and F1-score. These metrics were computed using the classification_report function from the sklearn.metrics library. Evaluation was performed on the validation set after each epoch to monitor the model’s progress and ensure effective learning.
4.5 Saving and Inference
After fine-tuning, the best-performing model was saved to disk for future use. The model and tokenizer were saved. For inference, we loaded the model using the Hugging Face pipeline API, simplifying predictions on new text samples. The model was tested on a sample Hausa text, and the predicted emotion was displayed.
4.6 Predictions on Test Set
To evaluate the model’s performance on unseen data and submit to SemEval 2025 Task 11, we made predictions on the test set. We loaded the test set from a Comma-Separated Values (CSV) file, and the model predicted the dominant emotion for each text sample. We converted the predictions into one-hot encoded labels and saved them to a new CSV file. This file contained the original text samples along with the predicted emotion labels, allowing for further analysis and evaluation.
5 Results and Discussion
Table 2: Training and validation metrics for the fine-tuned AfriBERTa model. The graph presents the accuracy and F1-score curves for both the training and validation sets across all training epochs, demonstrating the model’s learning progress and generalization capability
. Accuracy Precision Recall F1-Score Train 0.7416 ± 0.02 0.7515 ± 0.02 0.7348 ± 0.02 0.7393 ± 0.02 Val 0.7400 ± 0.02 0.7500 ± 0.02 0.7300 ± 0.02 0.7350 ± 0.02
The fine-tuned AfriBERTa model demonstrated consistent improvement over the training epochs, as shown in Table 2. The model achieved its best performance in the fifth epoch, with a training accuracy of 74.16% and an F1-score of 73.93%. Precision and recall scores were also strong, at 75.15% and 73.48%, respectively. These results indicate that the model effectively learned to classify emotions in Hausa text, despite the challenges posed by class imbalance and noisy data.
Figure 3: Comparison of Training and Validation Loss Across Epochs: The plot illustrates the progression of training and validation loss over five epochs. The training loss steadily decreases, indicating that the model is learning from the data. However, the validation loss initially decreases but later fluctuates, suggesting potential overfitting or variations in generalization performance.
Additionally, the model showed steady progress, with the training loss decreasing from 1.5168 in the first epoch to 0.3385 in the fifth epoch. The validation loss also decreased initially but stabilized around 0.8393 in the final epoch, suggesting that the model reached a point of convergence. The model’s performance on the validation set was consistent with the training results, achieving an accuracy of 74.00% and an F1-score of 73.50%, demonstrating its ability to generalize to unseen data.
The model performed particularly well on emotions like joy and sadness, which were more prevalent in the dataset. However, it struggled slightly with underrepresented emotions such as fear and disgust, likely due to their limited representation in the training data. This highlights the importance of addressing class imbalance in future work, potentially through techniques like data augmentation or weighted loss functions. Overall, the results underscore the effectiveness of AfriBERTa for emotion detection in low-resource languages and provide a strong baseline for future research in this domain.
6 Conclusion
In this work, we fine-tuned AfriBERTa for multi-label emotion detection in Hausa text, achieving an accuracy of 74.00% and an F1-score of 73.50%. The model performed well on prevalent emotions like joy and sadness but struggled with underrepresented emotions such as fear and disgust, highlighting the challenge of class imbalance. Our approach demonstrates the effectiveness of transformer-based models for low-resource language tasks. Future work could address class imbalance through data augmentation or weighted loss functions and extend this work to other African languages. This study provides a strong baseline for emotion detection in Hausa and contributes to the development of inclusive NLP systems.
7 Limitations
One major limitation of this work is the variability in Hausa dialects. The dataset primarily represents standard Hausa, which may not generalize well to regional dialects or informal variations commonly used in social media and conversational settings. This could lead to misclassification of emotions when processing text from underrepresented dialects.
Another limitation is the generality of the dataset. While the dataset is carefully curated, it may not fully capture the diversity of emotions expressed in different contexts, such as sarcasm, code-mixing with English or Arabic, and cultural-specific expressions. This limits the robustness of the model when applied to unseen, real-world data.
Additionally, computational constraints influenced the choice of AfriBERTa-small instead of larger transformer models. While this ensures efficiency, it may come at the cost of lower accuracy compared to models with higher capacity. The trade-off between computational efficiency and model performance is a key consideration for practical deployment.
Lastly, deep learning models often lack interpretability, making it difficult to explain why a particular emotion was predicted. This could pose challenges in high-stakes applications, such as mental health monitoring or crisis detection, where transparency is crucial.
Acknowledgements
The authors acknowledge Google DeepMind for a study grant.